Heya, im working on problem solving with systems of equations. i need help turning this word problem into a system of inequalities, so that i can graph them. heres the problem:
--You are recruiting students for a decorating committee for the Junior/Senior Prom. the committee can have 7 to 10 members. there must be more seniors then juniors on the committee.--
Im having a hard time even starting to convert this to a linear equality. Could you explain the steps for finding this out? Thanx
G

Answer by mukhopadhyay(490)

Assume that the number of Seniors members = x;
and the number of Junior members = y
Based on the problem,
x > y (because number of Seniors is more than the number of Juniors)
(x+y) >=7
(x+y) <=10
The linear inequalities are:
(x-y) >0
(x+y) >=7
(x+y) <=10
Hope this helps.